simran has brokerage account and intends to purchase 300 shares of TAta at r.s 40 per share. she has available capital is 8,000. she ask her broker for a broker’s call with the interest rate on the loan at 8% per year. the interest payment happens end of the year.

What is the percentage margin in your account when she purchased the 300 shares

If stock price falls to r.s 20 per share by the end of the year, what is the remaining margin in your account? (express margin in Ruppees)

Suppose price drops to R.S 20 per share at the end of the year, what will be her total rate of return of the investment?
Answer #1

Answer:

Purchase amount = 300 shares * 40 = 12,000

Capital available = 8,000

Capital requirement or margin required = 12,000 - 8000 = 4,000

Her margin = 8000/12000 =66.67%

Broker margin = 4000/12000=33.33%

Stock falls 20 per share

Then , the remaining margin = 300*20 = 6000

Simran margin account = 6000*66.67% = 4,000.

Total rate of return :-

Initial investment = (8000+4000)=12000

Total cost = 8000+4000+(4000*8%)

= 12,320

Profit or loss = final value at the end - cost

= (300*20) - 12,320

= 6000 - 12,320

= - 6320(loss)

Rate of return = - 6320/12,320*100 = - 51.30%